OpenStudy (haleyelizabeth2017):

A pulley of radius 15cm turns at 8 revolutions per second. What is the linear velocity of the belt driving the pulley in meters per second?

OpenStudy (anonymous):

\[\huge v=r \omega\] \(\omega\) is the angular velocity v is the linear velocity r is the radius

OpenStudy (samsungfanboy):

@AlexandervonHumboldt2 help me T_T or.. us lol

OpenStudy (haleyelizabeth2017):

I know, I need to multiply 8rev/sec by 15cm

OpenStudy (haleyelizabeth2017):

But wouldn't we need to multiply it by the circumference rather than just the radius?

OpenStudy (anonymous):

It gave you the frequency. From there you can calculate \(\omega\) and then plug that into the first equation I put

OpenStudy (anonymous):

Convert from rev/s to rpm (rev/min): \[\frac{ 8 rev }{ \sec }*\frac{ 60 \sec }{ minute }=\frac{ 480 rev }{ minute }=480 rpm\] \[60~ \text{rpm}=1~\text{Hz}\]\[480~\text{rpm} = 8~\text{Hz}\] Then: \[\huge \omega = 2 \pi f\]Where f = 8 Hz

OpenStudy (haleyelizabeth2017):

Well, a friend of mine says to use the conversion method...\[\frac{8rev}{sec}\times\frac{2\pi 15cm}{1 rev}\times\frac{1m}{100cm}\]

OpenStudy (haleyelizabeth2017):

lol oh

OpenStudy (anonymous):

Read this to make sure you make proper conversion for frequency. In order to use that specific equation for angular velocity, it must be in Hz https://en.wikipedia.org/wiki/Revolutions_per_minute

OpenStudy (haleyelizabeth2017):

:/

OpenStudy (anonymous):

You have to look at it backwards. It tells you that the frequency (in order to use that equation for \(\omega\)) must be in Hz. Well what is the equivalent for Hz? Well, it tell us the conversion that 60 rpm = 1 Hz. We're initially given rps, but we can easily convert the seconds to minutes in order to get rpm!

OpenStudy (samsungfanboy):

"When the pulley turns once, the belt moves one circumference (2 x pi x 15 cm) x by 8 to get distance per second in cm divide by 100 to get it into m/s "

OpenStudy (haleyelizabeth2017):

wait, I got about 7.54m/s

OpenStudy (samsungfanboy):

(;

OpenStudy (samsungfanboy):

here are you're answer choices 9.95 m/s 3.93 m/s 11.16 m/s 7.54 m/s

OpenStudy (anonymous):

Yep! I got that too using the method I described :)

OpenStudy (haleyelizabeth2017):

I got that using the method I posted ;)

OpenStudy (anonymous):

Yup! Also I realized that rev/s is already Hz. No need to convert to rpm bahaha. But yeah, use the formulas!

OpenStudy (haleyelizabeth2017):

lol thank you

OpenStudy (anonymous):

X) you're welcome